# AUTHOR ONE et al

### v2

@@ 2.2.1 Larger Language Models
The first stage is the statistical language modeling (SLM) stage. In this phase, the representation of words relies on solo heat vectors characterized by a sparse set of vectors representing words  89. Although this approach fails to reflect logical relations in the representation of word vectors, it provides an intuitive and understandable way of characterization. As for the characterization of sentences, Markov’s assumption is used to reveal the logical connections between words in a sentence through conditional probabilities. Specifically, this model fixes the context length, which is also called n-gram language model, among which the bigram and trigram models are the most common  11, 45. This phase of language modeling is mainly used for information retrieval  47, 33, 100. However, the main challenge faced by statistical language models is the “curse of dimensionality”, i.e., as the vocabulary increases, the parameter space required by the model grows exponentially. To alleviate this problem, techniques such as smoothing strategies are widely used to mitigate the effects of data sparsity  15, 72.
@@ 2.2.2 LLM-based Agents
The fourth stage is the Large Language Model-based agents (LLM-based agents) stage. As mentioned earlier, Large Language Models have demonstrated significant performance and can be utilized as the brain of LLM-based agents 43. LLM-based agents extend perception and action capabilities through strategies such as multimodal perception 102, 37 and tool invocation 44, 34, enhance planning and reasoning capabilities through techniques such as thought chaining and problem decomposition  35, 101, and gain the ability to interact with the environment by learning from and responding to feedback  84, 36. LLM-based agents acquire generalization capabilities by learning a large training set, thus enabling free switching between different tasks, and have been applied to various real-world tasks in areas such as financial services, smart homes, education and training, healthcare, and intelligent customer service.
@@
Figure 4: Trends in the number of arXiv papers published (in six-month increments) containing the keywords "Large Language Model" (since June 2018) and "Agents" (since June 2018), respectively. The statistics were calculated by querying the keywords in the title or abstract by month, using exact matches. We set different x-axis ranges for the two keywords. We labeled the points corresponding to important milestones in the progress of LLM and Agents’ research.The number of papers containing "Large Language Model" in the title or abstract increased dramatically after the release of ChatGPT (Fig.a). The number of papers containing "A" in the title or abstract increased dramatically after the release of AuotoGPT and HuggingGPT increased (Fig.b).
Figure 5: Illustration.
@@ 3.1 Knowledge Source
Figure 6: Distributed Knowledge Source.
@@
Figure 7: Distributed Knowledge.
@@
Multi-round conversation datasets involve not only linguistic conversations, but alignment between LLM-based agents and human commands in a broad sense. At the Large Language Model level, in order to understand human intentions, a multi-round Q-A (query-feedback) dataset, i.e., a large amount of query and feedback data, is required 80, abbasiantaeb2024letesteban2024using. In this process, the design of prompt templates (prompt word engineering) is crucial, which needs to guide the model to output the desired results without updating the weights of the Large Language Model according to the actual needs of the specialized domain. At the level of LLM-based agents, in order to realize the automatic maintenance of bridge structures, multiple rounds of I-A (instruction-action) datasets, i.e., a large number of instruction and action datasets, are required 23. Currently, a variety of intelligent algorithms and advanced devices have been developed in the field of bridge operation and maintenance, and the multi-round I-A dataset can help LLM-based agents invoke these tools more effectively to accomplish the automatic maintenance of bridges. It is worth noting that each tool invocation cannot be supported by a large number of algorithms and hardware development techniques.
@@ 4.1 Large Language Model for Bridge Operations and Maintenance Domain
Currently, Large Language Models in vertical fields are booming. In China, a number of industries have launched customized large language models, such as “Hua Tuo” in the medical and health industry, “Xuan Yuan” in the financial field, “Han Fei” in the legal industry, “Pangu Media” in the media and cultural communication field, and “Red Rabbit” in the field of intelligent customer service and marketing. Through deep learning and natural language processing technologies, these models accurately match the needs of various industries, providing a full range of intelligent solutions from medical knowledge quiz to financial risk assessment, from legal document generation to intelligent customer service, and strongly promoting the digital transformation and intelligent upgrade of various industries.
@@
Figure 8: Distributed Knowledge.
@@ 4.3 Automatic Generation of Bridge Reports and Assisted Decision-Making
Currently, most knowledge management Q&A systems in the field of bridge maintenance are based on knowledge graphs 99, 50, 67. Knowledge graph-based Q&A systems need to extract structured knowledge from bridge design, construction, operation, management, inspection, and other data, a process that is time-consuming and labor-intensive, and difficult to handle unstructured or semi-structured data  16. In addition, structured knowledge expressions may lead to error transmission and accumulation, and the accuracy of knowledge extraction decreases dramatically when data complexity increases.
@@
Figure 10: Distributed Knowledge.
4.4 Tools Calling LLM-based Agents
@@ 6 Conclusion
This study aims to propose a LLM-based agents based on large-scale language models to solve the current development bottlenecks in the field of bridge operation and maintenance, and then promote the process of industry intelligence. First, this paper overviews the development history of the bridge operation and maintenance field and LLM-based agents, analyzes the development bottlenecks faced by the bridge operation and maintenance field, and points out that the advantages of LLM-based agents in understanding, generating, planning, decision-making, and action can effectively solve these bottlenecks. Next, this paper elaborates the main framework for constructing LLM-based agents based on large-scale language modeling in the field of bridge operation and maintenance, including data sources, knowledge ontology construction, and model evaluation. In addition, this paper explores the potential development prospects of LLM-based agents based on large-scale language modeling in the field of bridge operation and maintenance. Finally, this paper analyzes the development trends and limitations of combining LLM-based agents with the bridge operation and maintenance domain, such as professional barriers and engineering ethical issues.
Although the combination of large-scale language model-based agents and vertical fields is mainly concentrated in a few industries with faster development of LLM-based agents, we firmly believe that the research and development of large-scale language models and intelligent agents will change the development direction of various industries. The operation and maintenance of bridge engineering, compared with design and construction, can give fuller play to the advantages of LLM-based agents, and with the development of the field of artificial intelligence, LLM-based agents than will shape a new pattern of intelligent operation and maintenance of bridges.

### v4

@@ 1 Introduction
In the context of the current rapid development of the research field of Large Language Models and Agents, the emergence rate of new Large Models and Agents has increased significantly. The academic community has paid great attention to this, and the publication of related research results has shown a sharp increase. This shows that AI technology is gradually penetrating and transforming various industries. Therefore, it is necessary to explore and introduce the corresponding advanced technology in the field of bridge operation and maintenance in order to promote the innovative development of this field.
@@
Figure 1: Trends in the number of arXiv papers published containing the keywords "Large Language Model" and "Agents" title or abstract by month, using exact matches. We set different x-axis ranges for the two keywords. We labeled the points corresponding to important milestones in the progress of LLM and Agents’ research.The number of papers containing "Large Language Model" in the title or abstract increased dramatically after the release of ChatGPT (Fig.a). The number of papers containing "A" in the title or abstract increased dramatically after the release of AuotoGPT and HuggingGPT increased (Fig.b).
@@ 2.2.1 Larger Language Models
The first stage is the statistical language modeling (SLM) stage. In this phase, the representation of words relies on solo heat vectors characterized by a sparse set of vectors representing words  89. Although this approach fails to reflect logical relations in the representation of word vectors, it provides an intuitive and understandable way of characterization. As for the characterization of sentences, Markov’s assumption is used to reveal the logical connections between words in a sentence through conditional probabilities. Specifically, this model fixes the context length, which is also called n-gram language model, among which the bigram and trigram models are the most common  11, 45. This phase of language modeling is mainly used for information retrieval  47, 33, 101. However, the main challenge faced by statistical language models is the “curse of dimensionality”, i.e., as the vocabulary increases, the parameter space required by the model grows exponentially. To alleviate this problem, techniques such as smoothing strategies are widely used to mitigate the effects of data sparsity  15, 72.
@@ 2.2.2 LLM-based Agents
The fourth stage is the Large Language Model-based agents (LLM-based agents) stage. As mentioned earlier, Large Language Models have demonstrated significant performance and can be utilized as the brain of LLM-based agents 43. LLM-based agents extend perception and action capabilities through strategies such as multimodal perception 103, 37 and tool invocation 44, 34, enhance planning and reasoning capabilities through techniques such as thought chaining and problem decomposition  35, 102, and gain the ability to interact with the environment by learning from and responding to feedback  84, 36. LLM-based agents acquire generalization capabilities by learning a large training set, thus enabling free switching between different tasks, and have been applied to various real-world tasks in areas such as financial services, smart homes, education and training, healthcare, and intelligent customer service.
@@ 3.1 Knowledge Source
Figure 5: Distributed knowledge can be used to train word embedding models or as external information for retrieval.
@@
Figure 6: Knowledge graphs as the most typical structured knowledge.
@@
Multi-round conversation datasets involve not only linguistic conversations, but alignment between LLM-based agents and human commands in a broad sense. At the Large Language Model level, in order to understand human intentions, a multi-round Q-A (query-feedback) dataset, i.e., a large amount of query and feedback data, is required 80, 99, 1. In this process, the design of prompt templates (prompt word engineering) is crucial, which needs to guide the model to output the desired results without updating the weights of the Large Language Model according to the actual needs of the specialized domain. At the level of LLM-based agents, in order to realize the automatic maintenance of bridge structures, multiple rounds of I-A (instruction-action) datasets, i.e., a large number of instruction and action datasets, are required 23. Currently, a variety of intelligent algorithms and advanced devices have been developed in the field of bridge operation and maintenance, and the multi-round I-A dataset can help LLM-based agents invoke these tools more effectively to accomplish the automatic maintenance of bridges. It is worth noting that each tool invocation cannot be supported by a large number of algorithms and hardware development techniques.
@@ 3.3 Qualitative Result Evaluation
Figure 7: An example of LLM-based agents framework.The main function of LLM-based agents is to perceive the service state of the bridge, then after the "brain" to think and take action. The "brain" needs to have bridge operation and maintenance knowledge system, can make predictions, evaluation, decision-making on bridge service performance.
@@ 4.1 Large Language Model for Bridge Operations and Maintenance Domain
Currently, Large Language Models in vertical fields are booming. In China, a number of industries have launched customized large language models, such as “Hua Tuo” in the medical and health industry, “Xuan Yuan” in the financial field, “Han Fei” in the legal industry, and “Red Rabbit” in the field of intelligent customer service and marketing. Through deep learning and natural language processing technologies, these models accurately match the needs of various industries, providing a full range of intelligent solutions.
@@
Figure 8: By parsing the pictures, we can analyze the location, main features, causes, impacts and disposal measures of bridge diseases, which cannot be done by the generalized Large Language Models in the market. The recognition of bridge disease pictures by Large Language Model can improve the professionalism and accuracy of bridge operation and maintenance, and reduce labor costs.
@@ 4.2 Multimodal Knowledge Graph for Bridge Operations and Maintenance Domain
Figure 9: LLM-based agents to realize multimodal (text, picture, video, table) parsing of bridge operation and maintenance information need to establish a knowledge system in various aspects such as bridge structure, disease, and evaluation. Knowledge graph can well increase the understanding depth of the model.
@@ 4.3 Automatic Generation of Bridge Reports and Assisted Decision-Making
Currently, most knowledge management Q&A systems in the field of bridge maintenance are based on knowledge graphs 100, 50, 67. Knowledge graph-based Q&A systems need to extract structured knowledge from bridge design, construction, operation, management, inspection, and other data, a process that is time-consuming and labor-intensive, and difficult to handle unstructured or semi-structured data  16. In addition, structured knowledge expressions may lead to error transmission and accumulation, and the accuracy of knowledge extraction decreases dramatically when data complexity increases.
Figure 10: Firstly, data preparation and preprocessing are carried out, and the database data are cleaned, segmented and mapped into word vectors, which are stored in the vector database. When the user question is received, it is preprocessed and mapped into word vectors, and the similarity with the segment vectors in the database is calculated to filter out the relevant information. According to the designed prompt template, the interrogative sentence and related information are input into the Large Language Model to generate a professional answer or report. Finally, user feedback is collected and the process is optimized to improve answer quality and user satisfaction.
4.4 Tools Calling LLM-based Agents
@@
Figure 11: LLM-based agents first use the monitoring system and remote inspection system to sense the bridge status, then use the multimodal processing capability (text, table, picture, video) to analyze the data in depth through the prediction and evaluation module, and integrate the raw data and its analysis process into the display platform. Eventually, LLM-based agents output decision-making results covering monitoring reports, inspection reports, maintenance programs, and develop action plans to implement bridge operation and maintenance operations by autonomously controlling intelligent devices.The current technology is still difficult to realize the unmanned and automated control of intelligent devices, but this must be the future direction of technological development.
@@ 6 Conclusion
In the course of this study, we have successfully constructed a LLM-based agents, aiming to break through the development bottleneck in the field of bridge operation and maintenance and to promote the intelligent transformation of this industry. By comprehensively analyzing the current development status of the bridge O&M field and the advancement of LLM-based agents, we conclude that LLM-based agents have significant advantages in understanding, generating, planning, decision-making, and action, and are able to effectively respond to the challenges facing the bridge O&M field. The framework of LLM-based agents proposed in this study covers several key aspects such as data sources, knowledge ontology construction and model evaluation, which provides strong support for the intelligent development of the bridge O&M domain. Meanwhile, we also explore the application prospect of this LLM-based agents in the bridge O&M field, and believe that it has a broad development space. However, the application of LLM-based agents to the field of bridge operation and maintenance still faces problems such as professional barriers and engineering ethics, which are important directions that need to be focused on and solved in future research. Overall, this study provides new ideas and methods for the development of intelligence in the field of bridge operation and maintenance, which is expected to promote the progress of the whole industry.